OpenStudy (anonymous):

HEELPPPP 3(x-1)=2(x+8)

OpenStudy (mathmale):

Multiply each of the 2 terms of (x-1) by 3. Please show your work.

rebeccaxhawaii (rebeccaxhawaii):

(3*x)-(3*-1)=(2*x)+(2*8)

OpenStudy (shadowlegendx):

\[3(x-1)=2(x+8) \] Distribute 3 to x and -1 then Distribute 2 to x and 8 When I say distribute, it means to multiply the terms in the parentheses by the number we are distributing(the number infront of the parentheses)

OpenStudy (shadowlegendx):

After we correctly distribute, we get this. \[3x - 3 = 2x + 16 \]

OpenStudy (shadowlegendx):

Now we isolate x

OpenStudy (shadowlegendx):

Add 3 to both sides \[3x = 2x + 19\]

OpenStudy (shadowlegendx):

Subtract 2x from both sides \[x = 19\]

rebeccaxhawaii (rebeccaxhawaii):

@ShadowLegendX shaddy let the user get the answer.

OpenStudy (anonymous):

thank you so so so much